# AIMCoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning

### v1

@@ Abstract
Multimodal Chain-of-Thought (CoT) has emerged as a powerful technique for enhancing the vision-language reasoning with interleaved information. However, existing methods often rely on simplistic heuristics for constructing interleaved CoT, typically depending on attention maps, which our empirical analysis reveals can be unreliable. What’s more, the shortcomings of their passive and purposeless selection strategies and their arbitrary triggering mechanisms in capturing the model’s cognitive need for information are further amplified. In this paper, we propose AIMCoT, an Active Information-driven Multi-modal Chain-of-Thought framework that addresses these fundamental limitations. AIMCoT introduces three synergistic components: (1) Context-enhanced Attention-map Generation (CAG), which mitigates the text-vision granularity imbalance, thereby producing more reliable attention maps as a foundation. (2) Active Visual Probing (AVP), which replaces passive selection with a proactive, goal-oriented strategy grounded in information theory to select image regions that help answer the questions maximally. (3) Dynamic Attention-shifting Trigger (DAT), which intelligently determines the optimal moments to insert visual information by monitoring the model’s text-to-vision attention shifts. Extensive experiments on three challenging benchmarks demonstrate that AIMCoT significantly outperforms state-of-the-art methods across different settings. By actively foraging for information and dynamically structuring its reasoning process, AIMCoT represents a critical step towards more robust, effective, and human-like multimodal reasoning. Our code is available at https://anonymous.4open.science/r/AIMCoT.
@@ 1 Introduction
The advent of Chain-of-Thought (CoT) prompting has marked a significant milestone in the reasoning capabilities of Large Language Models (LLMs) Wei et al. (2022); Wang et al. (2022); Zhang et al. (2022); Suzgun et al. (2022); Li et al. (2025); Yao et al. (2023); Besta et al. (2024), enabling them to deconstruct complex problems into a series of intermediate, interpretable steps. This paradigm has been naturally extended to Vision-Language Models (VLMs), where early efforts Mitra et al. (2024); Zheng et al. (2023); Lei et al. (2024); Zhang et al. (2023) focused on generating text-only rationales to articulate the model’s reasoning process over visual inputs. A pivotal advancement in this domain was the introduction of Interleaved-modal Chain-of-Thought Gao et al. (2025), which pioneered the direct integration of visual patches into the reasoning chain. By pairing textual rationales with corresponding image regions, the interleaved CoT demonstrated a superior ability to ground language in visual evidence, setting a new standard for multimodal reasoning.
However, the efficacy of existing research Ge et al. (2025); Gao et al. (2025) is fundamentally predicated on the reliability of their underlying mechanisms for selecting and integrating visual information. These methods typically rely on a passive, attention-driven strategy: they select the Top-K regions with the highest attention scores and insert them at predefined moments, e.g., the appearance of a newline character, a practice with little theoretical or empirical justification for its timing. This reliance exposes critical vulnerabilities that limit their full potential. Our empirical analysis reveals that high-attention regions are often redundant or, more alarmingly, fail to capture the most crucial visual details, especially when a granularity mismatch exists between the textual query and the visual evidence. This raises three fundamental questions: (1) What information source can reliably identify truly salient visual regions? (2) How can we accurately select useful regions in a proactive and purposeful manner rather than passively relying on potentially unreliable attention scores? (3) When is the optimal moment to insert visual evidence into the reasoning process?
In this work, we postulate that a more robust multimodal reasoning framework requires a shift from passive attention-following to an active, information-seeking paradigm. We propose AIMCoT: Active Information-driven Multi-modal Chain-of-Thought, a novel framework that directly addresses the limitations of prior work, reframing the selection of visual evidence from a passive, attention-based retrieval task to an active, goal-oriented probing process. Inspired by the principles of information foraging Pirolli & Card (1999); Broadbent (2013), AIMCoT operates on the premise that the most valuable visual rationale is one that maximally reduces the model’s uncertainty about the subsequent step in its reasoning chain. Rather than simply asking "Where is the model looking?", we prompt the model to actively ask "Which piece of visual information will be most helpful for me to see right now?". This shift is realized through three synergistic components:
  1. 1.
    Context-enhanced Attention-map Generation (CAG), which mitigates the text-vision granularity disparity to improve the reliability of attention map for better identifying salient regions by generating a context-aware description of the image.
  2. 2.
    Active Visual Probing (AVP), which implements a proactive, goal-oriented selection strategy grounded in information theory, selecting a set of visual regions that provide the highest possible information gain for the task at hand.
  3. 3.
    Dynamic Attention-shifting Trigger (DAT), an intelligent triggering mechanism that carefully captures the critical moments when model’s cognitive focus shifts significantly from text to vision, and inserts visual information precisely.
Our contributions are as follows:
  • We introduce AIMCoT, a novel training-free framework that reframes the construction of multimodal CoT as an active information-foraging process, moving beyond the limitations of static, passive, and purposeless region selection.
  • We propose a system comprised of three complementary methods (CAG, AVP, DAT) that collectively enable VLMs to proactively forage for informative visual evidence and strategically integrate it into their reasoning process at the detected critical moments.
  • We present a comprehensive set of empirically-grounded motivations that inspire the designs within our AIMCoT framework. Furthermore, we provide substantial theoretical analysis and design targeted experiments to explore the properties of these designs, covering key aspects such as the deployability of AIMCoT, the interplay between the CAG and AVP modules, and the necessity of incorporating an exploratory candidate pool.
  • Through extensive experiments conducted on two backbones (Chameleon-7B and Qwen2-VL-7B) across challenging benchmarks including M3CoT, ScienceQA, and LLaVA-W, we demonstrate that AIMCoT significantly outperforms state-of-the-art baselines and advances the frontier of vision-language reasoning.
@@ 2 Related Work
The success of CoT prompting in LLMs has naturally extended to VLMs, aiming to make their reasoning processes explicit and interpretable. Early efforts focused on generating text-only rationales. MMCoT Zhang et al. (2023) first generates a rationale from the input and then uses this rationale along with the original multimodal data to infer the final answer. CCoT Mitra et al. (2024) prompts the VLM to generate an intermediate scene graph to structure its understanding. DDCoT Zheng et al. (2023) decomposes complex problems into simpler sub-questions and leverages external models to fill information gaps. Alternatively, SCAFFOLD Lei et al. (2024) overlays a coordinate grid on the image, enabling the VLM to reference spatial regions explicitly in its textual reasoning.
A pivotal advancement is the introduction of interleaved-modal CoT, which integrates visual evidence directly into the reasoning chain. The leading approach, ICoT Gao et al. (2025), selects the top-K regions from the VLM’s attention map and inserts them at predefined moments. MRFD Ge et al. (2025) meticulously selects the salient visual regions to foster more reliable reasoning, alleviating VLM hallucination. However, our analysis reveals that such passive, attention-driven strategies are fundamentally limited by the unreliability of raw attention maps, especially during text-vision granularity mismatches, and their arbitrary insertion points fail to capture the model’s dynamic cognitive needs. Our proposed AIMCoT addresses these critical vulnerabilities by shifting the paradigm from passive selection to active information foraging, employing goal-oriented visual probing and dynamic, attention-shift-triggered integration.
@@ 3 Motivation
3.1 Identifying the Reliability of the Attention Map
The utilization of the attention map in multimodal learning has been explored by recent research Gao et al. (2025); Ge et al. (2025); Xie et al. (2022); Wang et al. (2023); Liu et al. (2022), inspiring us to also take it as a source from which the salient visual regions are selected. However, considering their strong dependence on attention map, we question its reliability by posing two fundamental questions:
@@ 3.1 Identifying the Reliability of the Attention Map
    Do all significant regions on the attention map help the VLMs answer questions correctly?
@@
    Does the attention map comprehensively capture all visual regions that are instrumental to the VLM’s correct prediction?
Figure 1: The images of the 22nd question on LLaVA-W benchmark, which is a close-up photo of a meal at ICHIRAN. The left and right figures are respectively the original image and the first three sets of regions selected by the Top-K strategy (red, purple, and blue, respectively). A detailed explanation is shown in Appendix B.
To acquire a holistic comprehension of these two questions, our empirical analysis is conducted on the popular Visual Question Answering (VQA) benchmark, LLaVA-W Liu et al. (2024), employing ICoT Gao et al. (2025) as a baseline model, which selects the Top-K regions from the attention map to construct a text-vision interleaved CoT. Chameleon-7B Team (2024) serves as the backbone.
Table 1: Performance degradation of the baseline model (ICoT, 0-shot) when the Top $K_{mask}$ regions on the attention map are masked.
$K_{mask}$ 0 1 5 10 20
Degradation 0% 0.26% 1.43% 3.93% 2.44%
We quantitatively investigate the role of high-attention regions by masking the top-$K_{mask}$ regions identified by ICoT (0-shot). On LLaVA-W, this masking leads to a minor performance drop of just 3.93% (top 10) and 2.44% (top 20), as shown in Table 1. Notably, performance increases when the mask is expanded from the top 10 to the top 20 regions. These results strongly suggest that not all high-attention regions contribute significantly to the model’s prediction: some have negligible impact or even introduce detrimental signals.
In response to the second question, we manually inspect the content of high-attention regions to understand their role in VLM’s predictions. In a challenging LLaVA-W example (Figure 1), the answer lies in a small area, i.e., the inner rim of a ramen bowl, among rich visual information. The top two most attended image patch sets entirely miss this crucial detail, with only a negligible portion appearing in the third set. This indicates a potential misalignment between high-attention regions and key visual information, particularly when the text-vision granularity disparity is significant.
3.2 Foraging for the Most Information to Guide Region Selection
The analysis in Section 3.1 empirically reveals that the efficacy of the attention-driven Top-K strategy is notably constrained, particularly when faced with significant text-vision granularity disparities. Furthermore, fundamentally, since attention maps merely reflect token correlations, a static approach without an explicit goal like Top-K, which relies solely on attention scores, is inherently suboptimal. This insight directly motivates us to explore a selection method that is more proactive and purposeful.
Existing recognized research Pirolli & Card (1999); Broadbent (2013); Oaksford & Chater (1994); Friston (2010) supports that people, when possible, will maximize their rate of gaining valuable information, as it yields more useful information per unit cost. Inspired by this, we explore providing the VLM with image regions that yield the highest information gain, thereby maximally reducing the model’s uncertainty in answering a given question. Intuitively, the model’s uncertainty can be quantified by the entropy of its probability distribution over the vocabulary given the current context, while the information gain of an image region is measured based on model’s entropy when the region is included in the context. The definitions are meticulously detailed in Section 4.3.
Figure 2: The visualization of regions selected by the information gain-guided strategy.
To continue with the example in Section 3.1, Figure 2 illustrates the top three regions ranked by the information gain for VLM. Evidently, in contrast to the regions selected via Top-K shown in Figure 1, the information gain-based selection accurately guides the VLM to first focus on the inner rim of the bowl (red), where the critical information is contained. Although the VLM does not yield the final answer in this region, we note that this indicates a correct line of reasoning, as the ground truth is situated in a highly similar area (a nearby location also on the inner wall of the bowl). Subsequently, the region ranked third (blue) precisely encompasses a large portion of the restaurant’s name, which is just the answer to the question. This suggests that even for a challenging case where the text-vision granularity is highly disparate, information gain serves as a better foundation for region selection.
3.3 Inserting Visual Information at Key Moments
Although existing research Gao et al. (2025) has attempted to construct interleaved CoT, they often ignore a reasonable triggering mechanism to capture the critical moment of inserting visual information. For example, ICoT Gao et al. (2025) simply uses a newline character as a trigger signal. This motivates an in-depth investigation into reliable indicators for identifying these key moments for inserting visual content.
To this end, we conduct a comparative case study on the LLaVA-W benchmark, with ROUGE-L as the evaluation metric. We collect all predictions generated by the baseline model (ICoT) and partition them into high- and low-scoring groups based on ROUGE-L scores. Given that the model’s attention map has served as an important foundation for existing works Gao et al. (2025); Ge et al. (2025); Xie et al. (2022); Wang et al. (2023); Liu et al. (2022), which is not only readily accessible but also effectively reflects the model’s aggregate attention to the textual and visual parts of the input, we are inspired to monitor the attention shifts between these two modalities throughout the prediction process. The experiments are detailed in Appendix F, from which two key observations are revealed: (1) Correlation analysis: inserting visual data precisely when the model’s attention pivots towards the visual modality is strongly correlated with higher scores and (2) Group analysis: this phenomenon further serves as a crucial characteristic that distinguishes high-scoring from low-scoring outputs.
4 AIMCoT
In this section, we begin by briefly reviewing the background of multimodal learning. Then, we detail the proposed methods motivated by the following key insights derived from Section 3:
  1. (1)
    First, the high-scoring attention regions are not always beneficial for question-answering, and crucial visual evidence can be missed particularly in cases of text-vision granularity mismatch.
  2. (2)
    Second, for a given set of candidate regions, selection based on information gain significantly outperforms the conventional attention-driven Top-K method.
  3. (3)
    Finally, capturing the critical moments to insert visual information improves the construction of multimodal CoT, and text-to-vision attention shifts serve as an important indicator.
Accordingly, we propose AIMCoT, which encompasses three key methods: (1) Context-enhanced Attention-map Generation (CAG), which generates a fine-grained description for the image to alleviate text-vision disparity. (2) Active Visual Probing (AVP), which proactively and purposefully selects regions that are most helpful for answering the question. (3) Dynamic Attention-shifting Trigger (DAT), which triggers vision insertion into CoT when the model’s cognitive focus is significantly shifted from text to vision.

### v2

# AIM-CoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning

@@ Abstract
Interleaved-Modal Chain-of-Thought (I-MCoT) advances vision-language reasoning, such as Visual Question Answering (VQA). This paradigm integrates specially selected visual evidence from the input image into the context of Vision-Language Models (VLMs), enabling them to ground their reasoning logic in these details. Accordingly, the efficacy of an I-MCoT framework relies on identifying what to see (evidence selection) and when to see it (triggering of insertions). However, existing methods fall short in both aspects. First, for selection, they rely on attention signals, which are unreliable—particularly under severe granularity imbalance between the brief textual query and the informative image. Second, for triggering, they adopt static triggers, which fail to capture the VLMs’ dynamic needs for visual evidence. To this end, we propose a novel I-MCoT framework, Active Information-driven Multi-modal Chain-of-Thought (AIM-CoT), which aims to improve both evidence selection and insertion triggering via: (1) Context-enhanced Attention-map Generation (CAG) to mitigate granularity imbalance via textual context enhancement; (2) Active Visual Probing (AVP) to proactively select the most informative evidence via an information foraging process; and (3) Dynamic Attention-shift Trigger (DAT) to precisely activate insertions when VLM’s attention shifts from text to visual context. Experiments across three benchmarks and four backbones demonstrate AIM-CoT’s consistent superiority. Our code is available at https://anonymous.4open.science/r/AIMCoT.
AIM-CoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning
Xiping Li1, Jianghong Ma2thanks: Corresponding Author 1The Chinese University of Hong Kong 2Harbin Institute of Technology (Shenzhen) lihsiping@gmail.com, majianghong@hit.edu.cn
@@ 1 Introduction
Chain-of-Thought (CoT) prompting, initially established for Large Language Models (LLMs) Wei et al. (2022); Wang et al. ; Zhang et al. (2023); Suzgun et al. (2023); Li et al. (2025); Wang et al. (2023); Diao et al. (2024); Chu et al. (2024), has been naturally adapted to Vision-Language Models (VLMs) Zhang et al. (2025); Chen et al. (2024b); Cheng et al. (2025); Xu et al. (2024), empowering them to tackle vision-related tasks via a series of intermediate reasoning steps. Visual Question Answering (VQA) serves as a representative scenario targeted in this study. In a typical VQA scenario, the VLM is presented with a multimodal question comprising a textual query and an associated image, and is then prompted to answer the question.
Early efforts in this domain Mitra et al. (2024); Zheng et al. (2023); Lei et al. (2025); Zhang et al. (2024) focused on generating text-only CoT. A pivotal advancement is the paradigm of Interleaved-Modal Chain-of-Thought (I-MCoT) Gao et al. (2025). Unlike text-only methods, this paradigm aims to provide fine-grained visual evidence for model reasoning. Specifically, it first selects the salient sub-regions from the input image, and then inserts them as visual tokens into the context of reasoning chain.
However, to construct an effective I-MCoT framework, two critical questions should be well addressed: (1) What to see? (Selection: identifying the specific image regions that support the current reasoning step) and (2) When to see it? (Triggering: determining the precise moment visual insertion is needed).
Existing research Gao et al. (2025) falls short in these two aspects due to its reliance on passive and static heuristics. Regarding selection, it adopts attention-based selection, which relies on the quality of cross-attention maps of VLM layers. Specifically, the visual regions that receive the most attention from tokens in the text context are selected, assumed to be most helpful for the subsequent generation. However, as revealed in Section 3.1, the raw VLM attention fails to accurately identify the salient regions, particularly when there is a granularity imbalance. Specifically, in the input context, the rich details in the image overwhelm the brief textual query. Therefore, with limited semantic anchors, the query cannot effectively steer the cross-attention towards the crucial visual regions. Regarding triggering, prior works often insert visual information at fixed, predefined moments (e.g., upon generating a newline character Gao et al. (2025)). As analyzed in Section 3.3, such a static mechanism fails to align with the model’s dynamic cognitive need for fine-grained visual information.
To overcome these limitations, we propose a novel I-MCoT framework, Active Information-driven Multi-modal Chain-of-Thought (AIM-CoT), which aims to address both what to see and when to see it by shifting VLM reasoning from passive, static perception to active, dynamic exploration. Grounded in Information Foraging Theory Pirolli and Card (1999); Broadbent (2013), AIM-CoT detects the VLM’s momentary need for visual cues and proactively forages for the most informative evidence. Integrated into the context window, this evidence improves subsequent generation. AIM-CoT realizes this shift through three synergistic components: (1) Context-enhanced Attention-map Generation (CAG): To refine the VLM’s cross-modal attention distribution, CAG elicits a query-conditioned description of the image and appends it to the textual context. Rather than serving as an auxiliary caption, CAG provides semantic anchors that align textual and visual granularities. (2) Active Visual Probing (AVP): To address the “What to see” problem, AVP quantifies the information gain of candidate regions, integrating those that provide the highest utility into the context for subsequent reasoning and generation. (3) Dynamic Attention-shift Trigger (DAT): To address the “When to see it” problem, DAT dynamically and precisely triggers visual insertion when a significant shift from textual to visual focus is detected, indicating the VLM’s cognitive demand for visual evidence.
Our contributions are summarized as follows:
  • We introduce AIM-CoT, a unified system (CAG, AVP, DAT) that enables VLMs to proactively forage for informative visual evidence and dynamically capture the critical triggering moments.
  • Our analyses demonstrate that AIM-CoT (1) selects truly salient visual evidence via information gain, (2) withstands noisy attention and descriptions, and (3) remains deployment-friendly.
  • Experiments conducted on four backbones across three VQA benchmarks demonstrate the consistently superior performance of AIM-CoT over state-of-the-art baselines.
@@ 2 Related Work
Multimodal CoT has been widely adopted in VLM research. Early text-only efforts enhance reasoning by decomposing questions Zheng et al. (2023), generating intermediate scene graphs Mitra et al. (2024), or overlaying coordinate grids for spatial referencing Lei et al. (2025).
A pivotal advancement is the I-MCoT paradigm, which integrates visual evidence directly into reasoning chains to improve subsequent reasoning/generation. The leading approach, ICoT Gao et al. (2025), selects the visual patches receiving the highest attention scores from the text tokens in the context. Then, this visual evidence is integrated when the VLM outputs a newline token.
However, our analysis in Section 3 highlights the limitations of existing research. Therefore, while we adopt the established I-MCoT paradigm (similar to ICoT Gao et al. (2025)), our contributions lie in shifting from the passive selection and static trigger to the proactive information-foraging selection and dynamic attention-shifting trigger.
@@ 3 Motivation
We focus on two pivotal questions for I-MCoT: What and When to see. First, we expose the limitations of passive attention-based selection in Section 3.1. To bridge this, we explore an active paradigm shift, investigating the potential of information gain for selection (Section 3.2) and dynamic attention shifts for triggering (Section 3.3).
3.1 Moving Beyond Attention Maps: The Reliability Gap in Passive Visual Selection
Existing I-MCoT methods adopt attention-based visual selection. We examine the reliability of attention for visual selection by asking:
@@ 3.1 Moving Beyond Attention Maps: The Reliability Gap in Passive Visual Selection
    Sufficiency: Do highly attended patches always benefit question answering?
@@
    Necessity: Do truly crucial patches always receive high attention?
To investigate these two questions, we randomly sample 500 instances from M3CoT and 500 from ScienceQA, respectively.
Regarding sufficiency, we mask the top-$K_{mask}$ most attended regions and evaluate the performance drop. Notably, the masking is only for evaluating the importance of these regions as visual evidence, not for AIM-CoT itself. As shown in Table 8 (Appendix C.1), masking highly attended regions leads to only minor degradation. Although information redundancy may explain this, our upcoming necessity check suggests that attention peaks rarely coincide with truly salient regions, failing to be sufficient signals for evidence selection.
Regarding necessity, we examine how well the most attended regions $R_{attn}$ cover the truly crucial regions $R_{true}$ in Section 5.5, which is quantified by Intersection over Union (IoU). Across 1,000 instances, over 75% have an IoU even below 50%. This shows that attention peaks rarely align with truly critical visual evidence, i.e., crucial regions do not necessarily emerge as the most attended ones. Moreover, Appendix C.2 suggests that alleviating the text-vision granularity imbalance is an effective way to improve this alignment.
In conclusion, attention alone is unreliable for evidence selection. Therefore, we do not treat attention as a selection rule. However, improving attention distribution remains crucial, as it (1) underlies autoregressive generation in VLMs and (2) provides candidate regions for subsequent information-driven selection. Accordingly, we introduce CAG, which assists VLMs to attend to critical visual content precisely by enriching brief textual queries with semantic anchors (Section 4.2).
3.2 Identifying What to See: From Semantic Alignment to Information Gain
The empirical limitations observed in Section 3.1 can be attributed to a fundamental misalignment: cross-attention maps primarily capture the semantic correlations among tokens, whereas the ultimate goal of the selection phase in I-MCoT is to identify visual evidence that provides rich visual information for subsequent reasoning and generation.
This insight motivates a shift from attention-based selection to an information-centric approach. However, two fundamental questions arise regarding the design of such a mechanism: (1) Metric: What constitutes a theoretical measure of “information” in the context of VLM reasoning? (2) Process: How should the selection mechanism mimic the cognitive process of information gathering, as opposed to static Top-K selection?
To answer these, we adopt Information Foraging Theory (IFT) Pirolli and Card (1999); Broadbent (2013); Oaksford and Chater (1994); Friston (2010) as a guiding principle for our framework design. First, regarding the metric, IFT suggests that information is valuable only when it reduces the agent’s uncertainty. This motivates us to ground our selection criterion in uncertainty reduction (i.e., information gain), which can be quantified by the entropy of the VLM’s predictive distribution over its vocabulary. Second, regarding the process, IFT characterizes foraging as a sequential trajectory where each step depends on previous knowledge. This suggests that the optimal strategy is not a one-off Top-K selection, but a dynamic, iterative process where the model updates its belief state after every glimpse, i.e., the selected regions are inserted into VLM’s context in the form of visual tokens.
(a) Regions selected by attention-based strategy.
(b) Regions selected by information-driven strategy.
Figure 1: A comparison of regions selected by different strategies. Details are provided in Appendix D.1.
These two theoretical insights directly shape the architecture of our AVP module (Section 4.3). Figure 1 compares AVP-selected regions with those from attention-based method, which is further detailed in Appendix D.1.
3.3 Identifying When to See: From Static Heuristics to Dynamic Attention Shifts
The fundamental goal of I-MCoT is to ground the VLM’s reasoning in inserted fine-grained visual evidence. Therefore, the insertions should align with the model’s fluctuating need for visual evidence across reasoning steps. However, existing methods rely on static triggers like newlines Gao et al. (2025), failing to capture critical moments when the VLM actively seeks visual evidence to support subsequent generation.
To address this, we propose that the attention shift from textual to visual contexts (i.e., the text-to-vision attention shift) serves as a superior, dynamic indicator. Intuitively, a significant pivot in attention towards the visual modality suggests the VLM’s cognitive demand for visual grounding. Unlike static triggers, monitoring these shifts allows for determining the precise “when” based on the model’s real-time internal state.
We empirically validate this intuition through an in-depth analysis of the baseline model on the LLaVA-W benchmark (detailed in Appendix F). Our investigation correlates the timing of visual insertions with model performance, revealing two key observations: (1) Correlation analysis: Synchronizing visual insertions with salient text-to-vision attention shifts strongly correlates with better generation quality. (2) Group analysis: This dynamic shift pattern distinguishes high-quality responses from low-quality ones.
In conclusion, although attention is unreliable as a decision rule for selection, its shift is still a reliable diagnostic signal revealing when the model is seeking visual information. Motivated by this, we present DAT, an attention-shift trigger that dynamically activates visual insertions (Section 4.4).
Figure 2: An overview of the AIM-CoT framework. CAG initially enhances the query to refine attention distributions, thereby supporting reasoning and supplying reliable candidates for AVP. During reasoning, AVP executes visual insertion by iteratively selecting the most informative regions as evidence conditioned on the current context. This insertion is triggered once DAT (the trigger) detects a significant text-to-vision shift in the VLM’s internal attention state. See Appendix D.2 for a concise reading guide.
4 AIM-CoT
In this section, we begin by briefly reviewing the background of vision-language reasoning. Then, based on the motivations detailed in Section 3, we present AIM-CoT. As a training-free framework, AIM-CoT models the interleaved-modal reasoning as an information-foraging process with three synergistic components: (1) CAG, which pre-processes the input to generate a fine-grained description, mitigating text-vision granularity imbalance. (2) AVP, which proactively selects regions that maximize information gain whenever activated. (3) DAT, which triggers AVP to insert visual evidence into the CoT precisely when the model’s cognitive focus shifts from text to vision. A visualization of AIM-CoT is shown in Figure 2.
It is important to distinguish AIM-CoT from attribution methods that rely on masking or deleting parts of the input image. AIM-CoT operates on a frozen VLM and adopts a “Trigger-Select-Insert” paradigm: although sub-regions are integrated as visual evidence, the input context, including the original image is fully preserved. This ensures the VLM operates on a coherent, augmented context rather than a disrupted one Khorram et al. (2021).